OpenStudy (anonymous):

Find x. Round to the nearest tenth if necessary. Assume that segments that appear to be tangent are tangent. x(x+5)=2(2+10)

OpenStudy (misty1212):

HI!

OpenStudy (misty1212):

not clear what that tangent part means, but if it is \[x(x+5)=2(x+10)\] then when you multiply out you get \[x^2+5x=2x+20\]which is a quadratic

OpenStudy (misty1212):

set equal to zero \[x^2+3x-20=0\] then the quadratic formula

OpenStudy (anonymous):

then what would I do after that ? @misty1212

OpenStudy (misty1212):

do you know the quadratic formula?

OpenStudy (misty1212):

oh dang i made a mistake!!

OpenStudy (misty1212):

is this the original question \[x(x+5)=2(2+10)\]??

OpenStudy (misty1212):

if so, then it is just \[x^2+5x=24\]so \[x^2+5x-24=0\] a totally different equation

OpenStudy (misty1212):

this one is much easier factor as \[(x-3)(x+8)=0\] and now you can find the zeros easily

OpenStudy (anonymous):

yes! thank you @misty1212